OpenStudy (anonymous):

solve the system of equations. x-y+z=8 x+y+z=6 x+y-z=-12

OpenStudy (anonymous):

i dont understand. i am trying to see what u mean.

OpenStudy (anonymous):

x-y+z=8 x+y+z=6 x+y-z=-12 Add first and last 2 x = -4 x=-2

OpenStudy (anonymous):

x-y+z=8 x+y+z=6 x+y-z=-12 and second and third 2x + 2 y=-6 2(-2) + 2 y = -6 2 y=-2 y=-1 Find z